# Quantum theory of fractional topological pumping of lattice solitons

### v2

@@ Abstract
One of the hallmarks of topological systems is the robust quantization of particle transport. It is the origin of the integer-valued quantum Hall conductivity and a potential tool for quantum information technology. Recent experiments on topological pumps constructed by using arrays of photonic waveguides and described by the Andre-Aubry-Harper (AAH) model, have demonstrated both integer and fractional transport of lattice solitons. In these systems, a background medium mediates interactions between photons via a Kerr nonlinearity and leads to the formation of self-bound multi-photon states. Upon increasing the interaction strength a sequence of transitions was observed from a phase with integer transport in a pump cycle through different phases of fractional transport to a phase with no transport. We here present a quantum description of topological pumps of self-bound many-particle states in terms of an effective Hamiltonian of their center-of-mass (COM) motion, which allows to classify topological phases in terms of generalized symmetries. We identify a topological invariant, an effective single-particle Chern number, which fully governs the soliton transport. Increasing the interaction strength in the AAH model leads to a successive merging of COM bands, which is the origin of the observed sequence of topological phase transitions and also the potential breakdown of topological quantization for some interaction strength.
@@ I Introduction
Topological quantum systems have been intensively studied since the discovery of the quantum Hall effect [1]. One of the simplest examples for such systems is a Thouless pump [2], which displays a quantized particle transport in an insulating bulk state of a 1D lattice upon cyclic adiabatic changes of system parameters. The transport is governed by an integer topological invariant, equivalent to a Chern number. The quantization of transport not only applies to a fully filled fermion band, it is also observable in the center-of-mass motion of a single particle equally distributed over all momentum states (see e.g. [3] for a detailed overview.) A major problem in the single particle case is the fast dispersion of the wavefunction. A possible solution for this has been utilized in [4, 5, 6] using bound many-particle objects: lattice solitons. They show quantized transport in a topological pump while being almost nondispersive due to their large mass. The notion solitons is used here colloquially as the self-bound many-particle states may not fulfill all properties of true solitons [7, 8, 9]. In the experiments of [6], laser pulses have been injected into spatially modulated waveguide arrays simulating a time-dependent Aubry-André-Harper (AAH) Hamiltonian with a Kerr nonlinearity mediating interactions [10, 11, 12]. Increasing the light intensity, solitons form, for which integer transport in a full pump cycle was observed. Above a certain power threshold, all transport is halted. In subsequent work [13], an interaction controlled transition between phases with integer and fractional transport was demonstrated (see Fig.1).
@@
These observations triggered an extended body of theoretical work [14, 15, 16, 17] based on the discrete nonlinear Schrödinger equation (DNLSE) [18]. While the DNLSE accurately reproduces the observed shift in the COM, the underlying mean-field description fails to provide an explanation for the quantized transport and of the origin of the topological phase transitions. Using perturbation theory it has been argued in [19, 14] that the transport is governed by the Chern number of the single-particle band from which the soliton bifurcates. However, already weak interactions destroy a uniform occupation of Bloch states, necessary for the one-to-one relation between particle transport and Chern-number of single-particle bands. Furthermore, as shown in Sec. II.2, there is a gradually growing admixture of higher single-particle bands when the interaction is increased, and the perturbative arguments fails entirely in the fractional case. Finally, cases of anomalous nonlinear Thouless pumping were found [20], and in [21] it was predicted that multi-component solitons can show fractional transport, despite the fact that the single-particle bands are all topologically trivial. This shows that the topological transport of solitons cannot be traced back to topological properties of the underlying single-particle Hamiltonian. Most recently a fermionic model with repulsive interactions has been investigated [22] where fractional transport emerges from coupling of single-particle bands by repulsive interactions of fermions. We here unravel the origin of the quantized soliton transport and provide an explanation for the observed phase transitions by developing a fully quantum description of topological soliton pumps. To be specific, we consider the AAH model as generic example; the approach applies, however, to all systems with self-bound many-particle states. We show that the topological contribution to their transport can be described in terms of an effective single-particle Chern number, provided the soliton band is gapped from all extended (not self-bound) many-body states. Introducing an effective single-particle Hamiltonian for solitons we show that in the AAH model for increasing interaction strength different soliton bands merge at some parameter values of the pump cycle. At this point, the transport is governed by a Wilson loop, giving rise to different phases with fractionally quantized average transport. Increasing the interaction further eventually mixes all soliton bands, and since the total Wilson loop of all soliton COM bands must vanish, the topological transport breaks down [6, 23]. Finally, we show that in some intermediate interaction regimes the energetically highest of a set of crossing soliton bands may be degenerate with extended states. In such a case the transport is no longer (fractionally) quantized and may take arbitrary values, explaining the fluctuating transport numerically predicted in [15]. Since the effective Hamiltonian of solitons is a single-particle Hamiltonian, its symmetries under time-reversal, charge-conjugation and chiral transformation provide a full classification of possible topological phases according to Ref. [24, 25].
@@ II.1 Aubry-André-Harper-model
$\enspace=-\sum_{l}\left[\bigl{(}J_{l}(t)\hat{a}^{\dagger}_{l}\hat{a}_{l+1}+h.c.\bigr{)}+\epsilon_{l}(t)\hat{a}_{l}^{\dagger}\hat{a}_{l}+\frac{U}{2}\hat{n}_{l}(\hat{n}_{l}-1)\right]$
@@
$J_{l}(t)=J\Bigl{(}1+\delta\cos{\bigl{(}\Omega t+\frac{2\pi lk}{p}\bigr{)}}\Bigr{)},$ (2)
with $0<\delta<1$, see Fig. 1. $p$, which we take as a prime number, whereas $k$ controls the phase shift between the hopping amplitudes. All on-site potentials $\epsilon_{l}$ in Eq. (1) are chosen to be 0.
@@
Attractive interactions $U$ lead to localized soliton states. These are states with a distribution of occupation numbers that decay with increasing distance to the center of mass with a localization length $\xi$, i.e. $\langle\hat{n}_{l+d}\hat{n}_{l}\rangle\,\sim\,\exp\{-|d|/\xi\}$ for $d\gg 1.$ We call them stable if they have an energy gap to all extended states with the same $K$. On the quantum level, a minimum value $U_{c}$ is required for a soliton to form, which tends to zero as $N\to\infty$ [26]. In a complex band structure multiple soliton solutions can exist, which for some parameter values may become degenerate. Furthermore, stable excited soliton solutions may not exist at all times, as they can become degenerate with extended states in some parts of the pump cycle. In this case we call these solitons partially stable.
@@ II.2 Mean-field approach: Discrete nonlinear Schrödinger equation
Numerical simulations of Eq. (3) have provided a good description of soliton energies and the observed soliton transport in parameter regimes where the soliton is stable. The semiclassical description, however, fails to explain the topological nature of the transport and the origin of topological phase transitions. In [19, 14] it was argued that for weak interactions the solitons can be represented in terms of the most localized Wannier states of only one single-particle Bloch band. The authors state that under this assumption the topological transport is described by the Chern number of the Bloch band the soliton bifurcates from.
In Fig. 2 we have plotted the integrated overlap of a soliton obtained from a self-consistent solution of Eq. (3) with the single particle Bloch wavefunctions for two lowest bands for the AAH model with phase offset $k=2$ and unit cell size $p=5$ as a function of interaction strength. The vertical spread of data points reflects different times in the pump cycle. One recognizes that while in the perturbative limit of very small absolute values of interaction strengths there is indeed a very large overlap with the lowest Bloch band, contributions from higher bands continuously grow with nonlinearity. In the regime of fractional transport, following the above argument, one would expect equal contributions from the lowest and the first excited band, which is clearly not the case. Moreover, if one calculates the overlap with the Bloch states in a momentum resolved way, one finds an inhomogeneous contribution from different lattice momenta, see [6]. Therefore, the resulting topological invariant is not just the average of the single-particle Chern-numbers.
Figure 2: Overlap of eigensolution of the DNLSE with single-particle Bloch bands for different interaction strength. Data points for the same interaction strength correspond to different times in the pumping cycle.
@@ III Center-of-mass transport and many-body Chern number
Since the semi-classical approaches to describe the transport of lattice solitons were shown to be incomplete, we will here derive a quantum mechanical description.Specifically we are interested in the transport of the COM of a soliton when ${\cal H}_{0}(t)$ is adiabatically varied over a period $T$, where the translational invariance is not changed. The instantaneous eigenstates of ${\cal H}(t)$ can be classified in terms of the conserved COM momentum $K$ and a band index $\mu$, and will be denoted as $|E_{\mu}(K)\rangle$ with energy $E_{\mu}(K)$.
@@
In the following we discuss the topological contributions to $d\langle\hat{X}\rangle/dt$ for degenerate and non-degenerate bands separately.
III.1 Non-degenerated soliton bands
Let us first consider a non-degenerate soliton band, i.e. for a given COM momentum $K$ there is only one eigenstate for each energy value in the whole pump cycle. In order to calculate the transport of a soliton from a single non-degenerate band, one can follow the well-known arguments for the transport of a single particle [2, 28]. To account for the topological transport starting in an instantaneous eigenstate $|E_{0}(K,t)\rangle$ at time $t$, one has to take into account non-adiabatic corrections. As the modulation of the Hamiltonian does not affect the translational invariance, non-adiabatic transitions only couple to states with the same COM momentum. Non-degenerate time-dependent perturbation theory yields
@@ III.1 Non-degenerated soliton bands
$\langle\hat{V}(q=0,t)\rangle=\frac{\partial E_{0}(K)}{\partial K}+$
@@
The first term in expression Eq. (6) describes the dynamical contribution to the transport. If we consider an initial state that has equal probability in all COM momenta $K$ in the Brillouin zone $\{-\pi,\pi\}$, the dynamical contribution vanishes. Then, the shift of the center of mass of the soliton in one period $T$ is given by the second term integrated over all COM momenta $K$
$\Delta\langle\hat{X}\rangle=i\int_{0}^{T}\!\!\!dt\int_{-\pi}^{\pi}\!\frac{dK}{2\pi}\,\left(\left\langle\frac{\partial E_{0}}{\partial t}\biggr{|}\frac{\partial E_{0}}{\partial K}\right\rangle-c.c.\right)=C.$ (9)
This is the effective single-particle Chern number of the soliton Bloch band. Note that we did not make any assumption about the many-body wavefunction other than its gapfulness. In particular, no assumption about its behavior in the relative coordinates of the particles was made. However, the assumption of gapfulness of the $N$-particle state with fixed COM momentum $K$ requires in general that the $N$ particles are bound to each other.
III.2 Degenerated soliton bands
In the degenerate case, i.e. if there are crossings of $m$ bands at some points in time, we need to apply degenerate time-dependent perturbation theory to an $m$ component vector $|\underline{\Psi}(K)\rangle=\bigl{(}|\Psi_{0}(K)\rangle,\dots,|\Psi_{m-1}(K)\bigr{)}^{\top}$ and can express the time-evolved state as [29]
@@ III.2 Degenerated soliton bands
is the Wilczek-Zee non-Abelian Berry phase, and non-adiabatic terms denotes the perturbative contributions due to non-adiabatic couplings to other, energetically separated states similar to the non-degenerate case. We note that since $\langle E_{l}|\partial_{t}E_{m}\rangle=\bigl{(}\langle E_{l}|\partial_{t}{\cal H}|E_{m}\rangle\bigr{)}/(E_{m}-E_{l})$ (compare Eq. (III.1) and (8)), the eigenstates of the matrix $\mathbf{A}_{K}$ coincide with the bare states $|E_{0}\rangle\dots|E_{m}\rangle$ far away from the crossing point in the adiabatic limit. At the crossing point the off-diagonal elements however diverge in general, which leads to a mixing.
Suppose the cyclic change of the Hamiltonian starts at a point where there is no degeneracy between COM bands and the system is prepared in one band, say $|E_{0}(K)\rangle$. Then, in the presence of isolated crossing points with other bands $|E_{\alpha}(K)\rangle$, a single cycle $t=0\to t=T$ will in general not return the initial state to itself, but multiple cycles are needed. Therefore, the topological transport is only integer quantized after multiple cycles, giving rise to a fractional average transport per cycle. In this case where there are crossings of soliton bands at some point in time, say of $|E_{0}(K,t)\rangle$ and $|E_{1}(K,t)\rangle$, the Chern number must be generalized to a Wilson loop
${C}_{n}=\frac{1}{2\pi}\int_{0}^{T}dt\,\partial_{t}\text{Im log det }\mathbf{P}(t)$ (16)
where $\mathbf{P}(t)={\cal T}\exp\left\{-i\int_{\textrm{BZ}}dK\,\mathbf{B}_{t}(K)\right\}$, and
$\mathbf{B}_{t}=\left(\begin{array}[]{cc}E_{0}&-i\langle E_{0}|\partial_{K}E_{1}\rangle\\ -i\langle E_{1}|\partial_{K}E_{0}\rangle&E_{1}\end{array}\right)$ (19)
is the Wilczek-Zee non-Abelian Berry phase [29] for fixed time, here for $n=2$ crossing bands. $C_{n}$ is an integer and the average topological transport per cycle in the $n$ bands is given by $C_{n}/n$.
The COM transport of a soliton is then fractional, provided it returns to its original band only after $n$ periods. We will show that this is the case for the Aubry-André-Harper model of Ref. [13].
IV Chern number and Wilson loop of solitons in the AAH model
While the soliton band structure can be well approximated by a self-consistent solution of the DNLSE, the many-body eigenstates and the Chern number, Eq. (9), or the Wilson loop, Eq. (16), must be obtained from solving the many-body Schrödinger equation, which constitutes a substantial challenge for more than a few particles. To tackle this problem we introduce the following basis of states with a fixed $K$ [30, 31]

### v3

@@ Abstract
One of the hallmarks of topological systems is the robust quantization of particle transport. It is the origin of the integer-valued quantum Hall conductivity and a potential tool for quantum information technology. Recent experiments on topological pumps constructed by using arrays of photonic waveguides and described by the (lattice-translational invariant) Aubry-André-Harper (AAH) model, have demonstrated both integer and fractional transport of lattice solitons. In these systems, a background medium mediates interactions between photons via a Kerr nonlinearity and leads to the formation of self-bound multi-photon states. Upon increasing the interaction strength a sequence of transitions was observed from a phase with integer transport in a pump cycle through different phases of fractional transport to a phase with no transport. We here present a quantum description of topological pumps of self-bound many-particle states in terms of an effective Hamiltonian of their center-of-mass (COM) motion, which allows to introduce an effective band structure $E_{\mu}(K)$ with $K$ being the COM momentum, and to classify topological phases in terms of generalized symmetries. We provide an explicit analytic expression of the effective Hamiltonian for few particles in the strong interaction limit and present numerical results in the more general case. We identify a topological invariant, an effective single-particle Chern number, which fully governs the soliton transport. Increasing the interaction strength in the AAH model leads to a successive merging of COM bands, which is the origin of the observed sequence of topological phase transitions and also the potential breakdown of topological quantization for some interaction strength.
@@ I Introduction
Topological quantum systems have been intensively studied since the discovery of the quantum Hall effect [18]. One of the simplest examples for such systems is a Thouless pump [41], which displays a quantized particle transport in an insulating bulk state of a 1D translationally invariant lattice upon cyclic adiabatic changes of system parameters. The transport is governed by an integer topological invariant, equivalent to a Chern number. The quantization of transport not only applies to a fully filled fermion band, it is also observable in the center-of-mass motion of a single particle equally distributed over all momentum states (see e.g. [5] for a detailed overview.) A major problem in the single particle case is the fast dispersion of the wavefunction. A possible solution for this has been utilized in [24, 29, 14] using bound many-particle objects: lattice solitons. They show quantized transport in a topological pump while being almost nondispersive due to their large mass. The notion solitons is used here colloquially as the self-bound many-particle states may not fulfill all properties of true solitons [19, 26, 3]. In the experiments of [14], laser pulses have been injected into spatially modulated waveguide arrays simulating a time-dependent Aubry-André-Harper (AAH) Hamiltonian with a Kerr nonlinearity mediating interactions [2, 8, 21]. Increasing the light intensity, solitons form, for which integer transport in a full pump cycle was observed. Above a certain power threshold, all transport is halted. In subsequent work [13], an interaction controlled transition between phases with integer and fractional transport was demonstrated (see Fig.1).
@@
These observations triggered an extended body of theoretical work [15, 48, 28, 42, 11, 35, 36] based on the discrete nonlinear Schrödinger equation (DNLSE) [22]. While the DNLSE accurately reproduces the observed shift in the COM and relates the transition points between phases of different fractional transport to bifurcation points of soliton solutions, the underlying mean-field description fails to provide an explanation for the quantized transport in terms of a topological invariant and of the origin of the topological phase transitions.
Using perturbation theory it has been argued in [15, 28] that for weak interactions the position of the soliton follows that of the Wannier centers of the single-particle band from which the soliton bifurcates, and thus is governed by the Chern number of this band. However, as shown in Sect. II.2, there is a gradually growing admixture of higher single-particle bands when the interaction is increased, and the perturbative arguments fail entirely in the fractional case. We note that there are approaches introducing nonlinear winding numbers [48, 36] or Chern numbers [35] to characterize topological properties of the mean-field solutions also in the regime of large nonlinearities. These approaches are however limited to systems with certain symmetries or capture only special nonlinear solutions which do not include self-bound solutions such as solitons or time-dependent Hamiltonians. Finally, cases of anomalous nonlinear Thouless pumping were found [38], and in [39] it was predicted that multi-component solitons can show fractional transport, despite the fact that the single-particle bands are all topologically trivial. This shows that the topological transport of solitons cannot be traced back to topological properties of the underlying single-particle Hamiltonian. Most recently a fermionic model with repulsive interactions has been investigated [16] where fractional transport emerges from coupling of single-particle bands by repulsive interactions of fermions.
We here unravel the origin of the quantized soliton transport and provide an explanation for the observed phase transitions by developing a fully quantum description of topological soliton pumps. To be specific, we consider the AAH model as generic example; the approach applies, however, to all systems with self-bound many-particle states in a translationally invariant lattice model. We show that the topological contribution to their transport can be described in terms of an effective single-particle Chern number, provided the soliton solution is gapped from all extended (not self-bound) many-body states. Introducing an effective single-particle Hamiltonian for solitons, which due to translational invariance can be characterized by an effective soliton band structure, we show that in the AAH model for increasing interaction strength different soliton bands merge at some parameter values of the pump cycle. At this point, the transport is governed by a Wilson loop, giving rise to different phases with fractionally quantized average transport. Increasing the interaction further eventually mixes all soliton bands, and since the total Wilson loop of all soliton COM bands must vanish, the topological transport collapses [14, 6]. For single-particle Bloch Hamiltonians also particles prepared in excited bands show quantized topological transport, determined by the corresponding Chern number. We show that an analogous behavior can be observed for solitons if the system is prepared in an excited soliton band that is stable and has a finite band gap to lower (and higher) lying solutions.
Finally, we show that in some intermediate interaction regimes the energetically highest of a set of crossing soliton bands may be degenerate with bands of extended states. In such a case the transport is no longer (fractionally) quantized and may take arbitrary values, explaining the fluctuating transport numerically predicted in [42]. Since the effective Hamiltonian of solitons is a single-particle Hamiltonian, its symmetries under time-reversal, charge-conjugation and chiral transformation provide a full classification of possible topological phases according to Ref. [1, 33].
@@ II.1 Aubry-André-Harper-model
$\kern 5.0pt=-\sum_{l}\left[\bigl(J_{l}(t)\hat{a}^{\dagger}_{l}\hat{a}_{l+1}+h.c.\bigr)+\epsilon_{l}(t)\hat{a}_{l}^{\dagger}\hat{a}_{l}+\frac{U}{2}\hat{n}_{l}(\hat{n}_{l}-1)\right]$
@@
$J_{l}(t)=J\Bigl(1+\delta\cos{\left(\Omega t+\frac{2\pi lk}{p}\right)}\Bigr),$ (2)
with $0<\delta<1$, see Fig. 1.
The spatial period $p$ is chosen to be a prime number so that it is never commensurable with $l$ and/or $k$ which would effectively reduce the spatial period. The phase-offset $k$, which is chosen to be smaller than $p/2$, determines the smallest possible fraction in the topological transport. If not explicitly stated otherwise, we choose a spatial period $p=5$ and a phase-offset $k=2$. This is the smallest possible unit-cell size where fractional transport appears in the AAH-chain. All on-site potentials $\epsilon_{l}$ in Eq. (1) are chosen to be 0.
@@
Attractive interactions $U$ lead to localized soliton states. These are states with a distribution of occupation numbers that decay with increasing distance to the center of mass with a localization length $\xi$, i.e. $\langle\hat{n}_{l+d}\hat{n}_{l}\rangle\,\sim\,\exp\{-|d|/\xi\}$ for $d\gg 1.$ We call them stable if they have an energy gap to all extended states with the same $K$. On the quantum level, a minimum value $U_{c}$ is required for a soliton to form, which tends to zero as $N\to\infty$ [30]. In a complex energy structure multiple soliton solutions can exist, which for some parameter values may become degenerate. Furthermore, stable excited soliton solutions may not exist at all times, as they can become degenerate with extended states in some parts of the pump cycle. In this case we call these solitons partially stable.
@@ II.2 Mean-field approach: Discrete nonlinear Schrödinger equation
Numerical simulations of Eq. (3) have provided a good description of soliton energies and the observed soliton transport in parameter regimes where the soliton is stable. The semiclassical description, however, fails to explain the topological nature of the transport and the origin of topological phase transitions. In [15, 28] it was argued that for weak interactions the solitons can be represented in terms of the most localized Wannier states of single-particle Bloch bands neglecting exchange terms between the latter. Under this assumption the topological transport is described by the Chern number of the Bloch band the soliton bifurcates from.
In Fig. 2 we have plotted the integrated overlap of a soliton obtained from a self-consistent solution of Eq. (3) with the single particle Bloch wavefunctions for the two lowest bands for the AAH model with phase offset $k=2$ and unit cell size $p=5$ as a function of interaction strength $U/J$. The vertical spread of data points reflects different times in the pump cycle. One recognizes that while in the perturbative limit of very small absolute values of interaction strengths there is indeed a close to unity overlap with the lowest Bloch band, contributions from higher bands continuously grow with increasing interaction strength. In particular, both in experiments and in numerical simulations of the DNLSE, integer transport is seen up to rather large interaction strength at which the admixture of higher Bloch bands is already quite substantial. Moreover, in the regime of fractional transport, following the above argument, one would expect equal contributions from the lowest and the first excited band, which is clearly not the case. Therefore, the resulting topological invariant is not just the average of the single-particle Chern-numbers.
Figure 2: Overlap of eigensolution of the DNLSE, Eq.  (3), for the AAH model with phase offset $k=2$ and unit-cell size $p=5$ with single-particle Bloch bands for different interaction strength $U/J$. Data points for the same interaction strength correspond to different times in the pumping cycle.
@@ III Center-of-mass transport and many-body Chern number
Since the semi-classical approaches to describe the transport of lattice solitons were shown to be incomplete, we will here derive a quantum mechanical description. Specifically, we are interested in the transport of the COM of a soliton when ${\cal H}_{0}(t)$ is adiabatically varied over a period $T$, where the translational invariance is not changed. The instantaneous eigenstates of ${\cal H}(t)$ can be classified in terms of the conserved COM momentum $K$ and a band index $\mu$, and will be denoted as $|E_{\mu}(K)\rangle$ with energy $E_{\mu}(K)$.
@@
In the following we discuss the topological contributions to $d\langle\hat{X}\rangle/dt$ for degenerate and non-degenerate solutions separately.
III.1 Non-degenerate soliton solutions
Let us first consider a non-degenerate soliton solution, i.e. for a given COM momentum $K$ there is only one eigenstate for each energy value in the whole pump cycle. In order to calculate the transport of a soliton from a single non-degenerate band, one can follow the well-known arguments for the transport of a single particle [41, 40]. To account for the topological transport starting in an instantaneous eigenstate $|E_{0}(K,t)\rangle$ at time $t$, one has to take into account non-adiabatic corrections. As the modulation of the Hamiltonian does not affect the translational invariance, non-adiabatic transitions only couple to states with the same COM momentum. Non-degenerate time-dependent perturbation theory yields
@@ III.1 Non-degenerate soliton solutions
$\langle\hat{V}(q=0,t)\rangle=\frac{\partial E_{0}(K)}{\partial K}$
@@
The first term in expression Eq. (6) is just the group velocity of the band and describes the dynamical contribution to the transport. If we consider an initial soliton state with coefficient $c(K)$ in momentum state $|E_{0}(K)\rangle$, which is symmetric with respect to $K=0$, the dynamical contribution vanishes. Then, the shift of the center of mass of the soliton in one period $T$ is given by the second term in Eq. (6) integrated over all COM momenta $K$
$\Delta\langle\hat{X}\rangle=\int_{0}^{T}\!\!\!dt\int_{-\pi}^{\pi}\!\!\!dK\,|c(K)|^{2}\,{\cal F}(K,t).$ (9)
Eq. (9) is a weighted integral over the Berry-curvature
${\cal F}(K,t)=i\left(\left\langle\frac{\partial E_{0}}{\partial t}\biggr|\frac{\partial E_{0}}{\partial K}\right\rangle-c.c.\right)$ (10)
in $(K,t)$ space. In the special case, where the COM wavefunction of the soliton is initially localized to a single unit cell, all coefficients are equal, $|c(K)|^{2}=1/{2\pi}$. In this case the transport is given by the effective single-particle Chern number of the soliton Bloch band, i.e. it is integer quantized:
$\Delta\langle\hat{X}\rangle=\int_{0}^{T}\!\!\!dt\int_{-\pi}^{\pi}\!\frac{dK}{2\pi}\,{\cal F}(K,t)=C.$ (11)
Note that we did not make any assumption about the many-body wavefunction other than its gapfulness. In particular, no assumption about its behavior in the relative coordinates of the particles was made. However, the assumption of gapfulness of the $N$-particle state with fixed COM momentum $K$ requires in general that the $N$ particles are bound to each other.
Finally we comment on the relation between our description of soliton transport and other topological invariants for interacting many-body systems, such as the Niu-Thouless (NT) invariant [31]. The NT invariant employs twisted boundaries and is an integral over a Berry curvature in $(\theta,t)$ space, with $\theta\in\{0,2\pi\}$ being the twist angle. In contrast to ${\cal F}(K,t)$, see Eq. (9), a Berry curvature in $(\theta,t)$ would not describe the transport of an individual soliton but applies only to an insulating many-body state, such as a band insulator of non-interacting fermions, where all momentum states are equally occupied.
III.2 Degenerate soliton solutions
In the degenerate case, i.e. if there are crossings of $m$ solutions at some points in time, we need to apply degenerate time-dependent perturbation theory to an $m$ component vector $|\underline{\Psi}(K)\rangle=\bigl(|\Psi_{0}(K)\rangle,\dots,|\Psi_{m-1}(K)\bigr)^{\top}$ and can express the time-evolved state as [46]
@@ III.2 Degenerate soliton solutions
is the Wilczek-Zee non-Abelian Berry connection, and ”non-adiabatic terms” denotes the perturbative contributions due to non-adiabatic couplings to other, energetically separated states similar to the non-degenerate case. We note that since $\langle E_{l}|\partial_{t}E_{m}\rangle=\bigl(\langle E_{l}|\partial_{t}{\cal H}|E_{m}\rangle\bigr)/(E_{m}-E_{l})$ (compare Eq. (III.1) and (8)), the eigenstates of the matrix $\mathbf{A}_{K}$ coincide with the bare states $|E_{0}\rangle\dots|E_{m}\rangle$ far away from the crossing point in the adiabatic limit. At the crossing point the off-diagonal elements however diverge in general, which leads to a mixing.
Suppose the cyclic change of the Hamiltonian starts at a point where there is no degeneracy between COM bands and the system is prepared in one solution, say $|E_{0}(K)\rangle$. Then, in the presence of isolated crossing points with other solutions $|E_{\alpha}(K)\rangle$, a single cycle $t=0\to t=T$ will in general not return the initial state to itself, but multiple cycles are needed. Therefore, the topological transport is only integer quantized after multiple cycles, giving rise to a fractional average transport per cycle. In this case where there are crossings of soliton solutions at some point in time, say of $|E_{0}(K,t)\rangle$ and $|E_{1}(K,t)\rangle$, the Chern number must be generalized to a Wilson loop
${C}_{n}=\frac{1}{2\pi}\int_{0}^{T}dt\,\partial_{t}\text{Im log det }\mathbf{W}(t)$ (18)
where $\mathbf{W}(t)={\cal T}\exp\left\{i\int_{\textrm{BZ}}dK\,\mathbf{B}_{t}(K)\right\}$, and
$\mathbf{B}_{t}=\left(\begin{array}[]{cc}\langle E_{0}|\partial_{K}E_{0}\rangle&i\langle E_{0}|\partial_{K}E_{1}\rangle\\ i\langle E_{1}|\partial_{K}E_{0}\rangle&\langle E_{1}|\partial_{K}E_{1}\rangle\end{array}\right)$ (21)
is the Wilczek-Zee non-Abelian Berry connection [46] for fixed time, here for $n=2$ crossing bands. $C_{n}$ is an integer and the average topological transport per cycle in the $n$ bands is given by $C_{n}/n$.
The COM transport of a soliton is then fractional, provided it returns to its original energy only after $n$ periods. We will show that this is the case for the Aubry-André-Harper model of Ref. [13].
IV Calculating Chern number and Wilson loop of solitons
While the soliton energy structure can be well approximated by a self-consistent solution of the DNLSE, the many-body eigenstates and the Chern number, Eq. (11), or the Wilson loop, Eq. (18), must be obtained from solving the many-body Schrödinger equation, which constitutes a substantial challenge for more than a few particles. To tackle this problem we introduce the following basis of states with a fixed $K$ [34, 17]